OpenStudy (shalonda):

whats -3 5/6- (-4 1/12)equals

OpenStudy (saifoo.khan):

\[-3\frac{5}{6}-(-4\frac{1}{12})\] Like this right?

OpenStudy (anonymous):

-49/12-23/6

OpenStudy (shalonda):

yes

OpenStudy (saifoo.khan):

\[\frac{1}{4}\]

OpenStudy (shalonda):

can you all plz explain

OpenStudy (saifoo.khan):

Sure, do u know the mixed fraction to improper frac? Like Joeway did?

OpenStudy (anonymous):

I got it wrong 25/3

OpenStudy (saifoo.khan):

But my answer is right.

OpenStudy (anonymous):

yes her answer is correct